# PALM: Pushing Adaptive Learning Rate Mechanisms for Continual Test-Time Adaptation

### v1

@@ 4.2 Parameter Sensitivity as an Indicator of Domain Shift
New studies have found that model parameters demonstrate levels of criticality [2, 50] across different positions. Moreover, not all parameters adapt uniformly, and removing some can lower test generalization errors [31, 34]. Inspired by this observation, we claim that parameters have different contributions during test-time adaptation. From Sec. 4.1, we adaptively select layers that are uncertain in their predictions. For parameters of these layers only, we compute their sensitivity i.e. an approximation of an error when it is removed.
In our work, motivated from [19, 30, 29], we compute the parameter sensitivity of a parameter $\theta_{n}^{t}$ based on the loss (Eq. 4) as follows,
$S_{n}^{t}=\mathcal{L}({\theta}^{t})-\mathcal{L}({\theta}^{t}|\theta_{n}^{t}=0)$ (7)
@@
Precisely, $S_{n}^{t}$ approximates the change in the loss landscape at $\theta^{t}$ when $\theta_{n}^{t}$ is removed. Sensitivity, computed as a "weighting factor" of model weights [30, 29, 23], determines the level of adaptation needed for model parameters.
We utilize the gradients of parameters from the selected layers, computed by Eq. 5 to compactly derive Eq. 7 from the first-order Taylor series expansion of $\mathcal{L}(\theta^{t})$ as,
$S_{n}^{t}=|\theta_{n}^{t}\nabla\mathcal{L}({\theta_{n}}^{t})|$ (8)
@@
Our rationale is as follows: parameters with low sensitivity are pushed to have higher learning rates owing to more adaptation. However, $S_{n}^{t}$ of $\theta_{n}^{t}$ can be highly uncertain and would have a large variation. We introduce a method to capture this "domain-level" uncertainty by computing a weighted moving average of the sensitivity $S_{n}^{t}$. This is formulated as,
$\hat{S}_{n}^{t}=\alpha S_{n}^{t}+(1-\alpha)\hat{S}_{n}^{(t-1)}$ (9)
@@
where $\hat{S}_{n}^{t}$ denotes the "domain-level" sensitivity and $\alpha$in [0,1] is a smoothing factor. Specifically, $\alpha$ decides the contribution of the current sensitivity $S_{n}^{t}$ to $\hat{S}_{n}^{t}$. Since CTTA happens at batch level, it is critical to achieve a good balance between $S_{n}^{t}$ and $\hat{S}_{n}^{t}$, with a slightly larger focus on $S_{n}^{t}$. As it can be seen, we do not directly accumulate the "domain-level" information, unlike LAW [33].
In addition, due to large distribution shifts happening because of the continuum of tasks, larger deviations of $S_{n}^{t}$ from $\hat{S}_{n}^{t}$ are bound to happen. To quantify these statistics, an absolute difference between $\hat{S}_{n}^{t}$ and $S_{n}^{t}$ can be computed as,
$\hat{D}_{n}^{t}=|S_{n}^{t}-\hat{S}_{n}^{t}|$ (10)
@@
We emphasize that since Eq. 9 leverages sensitivity from previously computed batches and domains, $\hat{D}_{n}^{t}$ would act as a good indicator of how uncertain $S_{n}^{t}$ is to the previous batches and domains, for $\theta_{n}$. This would take care of error accumulation too.
To automatically alter the learning rates of these parameters, the learning rate "importance" ($i_{n}^{t}$) is calculated as follows,
$i_{n}^{t}=(\hat{D}_{n}^{t})/(\hat{S}_{n}^{t})$ (11)
@@
Then, the final learning rate to update $\theta_{n}^{t}$ would be
$\hat{\kappa}_{n}^{t}=\kappa*i_{n}^{t}$ (12)
@@
where $\kappa$ is the base fixed learning rate. The intuition behind the formulation of Eq. 11 is as follows - smaller values of $\hat{S}_{n}^{t}$ indicate lower sensitivity and hence larger parameter updates are required via promoting the learning rate. At the same time, large uncertainties by $\hat{D}_{n}^{t}$ would also require larger updates. To avoid any overflow/underflow, we add $\epsilon$ (0<$\epsilon$$\ll$1) to the numerator and the denominator in Eq. 11.
@@ 4.3 Optimization Objective
Though our method is robust, it is necessary to avoid miscalibrated labels. TENT [44] minimized the Shannon entropy on the output logits, defined as,
$\mathcal{H}(\hat{y})=-\sum_{c}p(y_{c})log(p(y_{c}))$ (13)
where $p(y_{c})$ refers to the softmax probability of class c i.e. $p(y_{c})$ = softmax($h(x_{k})$). Following previous CTTA work [41], we use entropy minimization as the main loss objective for adaptation where Eq. 13 is minimized for samples below a certain threshold.
$\mathcal{L}^{entropy}=\mathds{1}_{\mathcal{H}(\hat{y})\leq\mathcal{H}_{0}}\mathcal{H}(\hat{y})$ (14)
where $\mathds{1}$ is the indicator function and $\mathcal{H}_{0}$ is a fixed threshold set to $0.4*lnC$, consistently following [41]. To be fair with prior works [45, 33], we use the consistency loss between the model predictions on $x_{k}$ and its augmented version $\hat{x}_{k}$ to regularize the model as follows,
$\mathcal{L}^{const}=-\sum_{c}p(y_{c})log(p(\hat{y}_{c}))$ (15)
where, $\hat{y}_{c}$ = $h(\hat{x}_{k})$. Our overall optimization objective is as follows,
@@
where $\lambda$ indicates a regularisation coefficient to balance the consistency loss. In summary, upon calculating the parameter-specific learning rates (Eq. 12), we optimize the continually fine-tuned model to minimize Eq. 16.
@@ 5 Experiments
We draw a contrast of our proposed method against other state-of-the-art works in a continual test-time adaptation (CTTA) setting. However, to have a broader study and analysis, we also report results in a gradual test-time adaptation (GTTA) setting, following [45, 33]. In the following subsections, we describe the implementation details of both the challenging adaptation settings. For additional results and analysis, please refer to the Supplementary Material.
@@ 5.1 Experimental Settings and Baselines
Datasets and tasks. Following the standard benchmarks in CTTA, set by CoTTA [45], we evaluate our proposed method on datasets - CIFAR-10C, CIFAR-100C, and ImageNet-C, based on image corruption schemes as set in [11]. Each dataset contains a set of 15 corruption styles (e.g. gaussian noise, shot noise, impulse noise, defocus blur) with 5 severity levels. In line with established practices outlined in [45, 6], we report the mean classification error as the evaluation metric.
CTTA implementation details. We evaluate our approach using frameworks from prior state-of-the-art methods [44, 45, 3, 41, 37, 33]. To maintain fairness regarding the usage of pre-trained vision models, we employ WideResNet-28 [48] for CIFAR-10C, ResNeXt-29 [46] for CIFAR-100C, and ResNet-50 [10] for ImageNet-C, all available on RobustBench [5]. We optimize using an Adam optimizer [17], setting base learning rates ($\kappa$) to 5e-4 for CIFAR-10C and CIFAR-100C, and 5e-5 for ImageNet-C. For balancing parameter sensitivities, we set $\alpha$ to 0.5, 0.9, and 0.5 respectively, with temperature coefficients T set to 50, 100, and 1000 respectively. We set $\eta$ to 1, 0.5, and 0.3, and $\lambda$ to 0.01 throughout. Batch sizes are set to 200, 200, and 64 for each dataset, following CoTTA [45], and results are reported for the highest severity level of 5 for each task.
GTTA implementation details. In the standard CTTA setup, each task receives images with the highest severity. However, in GTTA, following [45, 41, 33], for each task the severity levels are gradually changed as …3$\rightarrow$2$\rightarrow$1 $change]{task}$ 1$\rightarrow$2$\rightarrow$3$\rightarrow$4$\rightarrow$5$\rightarrow$4$\rightarrow$3$\rightarrow$2$\rightarrow$1 $change]{task}$ 1$\rightarrow$2$\rightarrow$3… We maintain consistency with the experimental details outlined in the CTTA setting.
Baselines. We compare our work against several CTTA baselines, including Source, BN Stats Adapt [37], TENT Continual [44], CoTTA [45], EcoTTA [41], Adacontrast [3], and LAW [33]. In addition, we also contrast our work against surgical fine-tuning [21]. For EcoTTA, we report the results on CIFAR-10C and ImageNet-C only since they do not use ResNeXt-29 [46] for CIFAR-100C. Following the findings of surgical fine-tuning [21], we continually fine-tune only the first convolutional block of the respective models, denoted as Surgical fine-tuning (continual). We have mentioned the training details in the Supplementary. For the reported results of LAW [33], we reproduce their results strictly following their official implementation  [^1] https://github.com/junia3/LayerwiseTTA/tree/main with parameters as mentioned in their paper.
Table 1: Mean errors (in %) on CIFAR-10C - We report the CTTA mean errors of the 15 corruptions at a severity level of 5. We report the error of a single update (online) only using WideResNet-28 [48] for all the reported baselines.
@@
25.89 18.1 22.71 12.38 25.3 13.18 10.79 13.5 13.15 12.21 8.53 11.85 17.98 12.02 15.47 15.54
Table 2: Mean errors (in %) on CIFAR-100C - We report the CTTA mean errors of the 15 corruptions at a severity level of 5. We report the error of a single update (online) only using ResNeXt-29 [46] for all the reported baselines.
@@
37.33 32.5 34.94 26.20 35.28 27.54 24.64 28.85 29.15 34.14 23.52 26.99 31.15 26.56 34.1 30.19
@@ 5.2 CTTA Classification Results
Results on CIFAR-10C and CIFAR-100C. We contrast our proposed method against several baselines including surgical fine-tuning [21]. Tables 1 and 2 report the CTTA mean classification errors across all the domains, for each of the methods. The results reveal that our method performs the best across all training setups and baselines for both datasets. In particular, we achieve large improvements of 4.86% and 5.16% for CIFAR-10C and 5.21% and 30.71% for CIFAR-100C, over BN Stats Adapt[3] and TENT Continual[44] respectively. Concerning recent work LAW [33] that uses an adaptive LR scheme, our method achieves an improvement of 1.67% and 1.56% on CIFAR-10C and CIFAR-100C, respectively. It’s noteworthy that TENT Continual initially excels in the early tasks but experiences error accumulation over time, leading to an increase in mean error. In the study by CoTTA [45], despite achieving good performance, the approach’s reliance on full model updates renders it computationally expensive and unsuitable for deployment. LAW [33] uses the uncertain pseudo labels, of each batch, to calculate the domain shift approximated via the FIM computation. The minor error that is introduced for each batch begins to propagate across all the domains. In addition, the imbalance caused by the "domain-level" FIM and the current batch’s FIM, results in a higher error rate for the initial tasks. In comparison, the computation of the model prediction uncertainty to select layers is useful in our work (Eq. 4). Another interesting observation is that with a drastic reduction in the number of trainable parameters, as in Surgical fine-tuning (continual), the mean errors in Table 2 is better than most CTTA benchmarks, if not the best. This affirms the findings in [21] that concentrated adaptation, for image corruptions, in the initial layers is sufficient.
Results on Imagenet-C In Table 3, we report the CTTA results on the ImageNet-C dataset, with ResNet-50 pre-trained on ImageNet. As shown, our method outperforms all the previous CTTA baselines and Surgical fine-tuning (continual). We observe an improvement of 1% over LAW [33]. In addition, we achieve 2.31%, 5.38%, and 30.71% improvements over CoTTA [45], Adacontrast [3], and TENT [44], respectively.
Table 3: Mean errors (in %) on ImageNet-C - We report the CTTA mean errors of the 15 corruptions at a severity level of 5. We report the error of a single update (online) only using ResNet-50 [10] for all the reported baselines.
@@
81.06 73.32 70.88 77.00 71.88 62.34 53.94 56.7 60.76 50.36 36.26 65.96 48.08 45.28 48.0 60.12
Table 4: Mean errors (in %) on CIFAR-10C, CIFAR-100C, and ImageNet-C - We report the gradual TTA (GTTA) mean errors averaged across all the severity levels (1normal-$\rightarrow$2normal-$\rightarrow$3normal-$\rightarrow$4normal-$\rightarrow$5).
@@
CIFAR-10C 24.7 13.7 20.4 10.9 12.1 11.33 10.87 9.56
CIFAR-100C 33.6 29.9 74.8 26.3 33.0 25.31 27.19 26.11
ImageNet-C 58.4 48.3 46.4 38.8 66.3 41.36 40.52 38.49

### v2

@@ 4.2 Parameter Sensitivity as a Domain Shift Indicator
New studies have found that model parameters demonstrate levels of criticality [2, 44] across different positions. Moreover, not all parameters adapt uniformly, and removing some can lower test generalization errors [22, 25]. Inspired by this observation, we assert that parameters contribute differently during CTTA. From the previous stage, we automatically select the layers based on the prediction uncertainty. For parameters of these selected layers only, we compute their sensitivity i.e., an approximation of an error when it is removed. Motivated by [14, 21, 20], we compute the sensitivity of a parameter $\theta_{j,n}^{t}$ based on the loss (Eq. 1) as follows,
$S_{j,n}^{t}=\mathcal{L}({\theta}^{t})-\mathcal{L}({\theta}^{t}|\theta_{j,n}^{t}=0)$ (3)
@@
Precisely, $S_{j,n}^{t}$ approximates the change in the loss landscape at $\theta^{t}$ when $\theta_{j,n}^{t}$ is removed. Sensitivity, computed as a “weighting factor" of model weights [21, 20, 16], determines the level of adaptation needed for model parameters. We utilize the gradients of parameters from the selected layers, computed by Eq. 2, to compactly derive Eq. 3 from the first-order Taylor series expansion of $\mathcal{L}(\theta^{t})$ [lee2018snip] to,
$S_{j,n}^{t}=|\theta_{j,n}^{t}\nabla_{\theta_{j,n}^{t}}\mathcal{L}({\theta^{t}})|$ (4)
@@
Our rationale is as follows: parameters with low sensitivity are pushed to have higher LRs owing to more adaptation. However, $S_{j,n}^{t}$ of $\theta_{j,n}^{t}$ can be highly uncertain and would have a large variation. We introduce a method to capture this “domain-level" uncertainty by computing a weighted moving average of the sensitivity $S_{j,n}^{t}$. This is formulated as,
$\hat{S}_{j,n}^{t}=\alpha S_{j,n}^{t}+(1-\alpha)\hat{S}_{j,n}^{(t-1)}$ (5)
@@
where $\hat{S}_{j,n}^{t}$ denotes the “domain-level" sensitivity, computed on-the-fly, and $\alpha$in [0,1] is a smoothing factor. Specifically, $\alpha$ decides the contribution of the current sensitivity $S_{j,n}^{t}$ to $\hat{S}_{j,n}^{t}$. Since CTTA happens at the batch level, it is critical to achieve a good balance between $S_{j,n}^{t}$ and $\hat{S}_{j,n}^{t}$, with a slightly larger focus on $S_{j,n}^{t}$. As can be seen, we do not directly accumulate the “domain-level" information, unlike LAW. Due to large distribution shifts happening because of the continuum of tasks, larger deviations of $S_{n}^{t}$ from $\hat{S}_{n}^{t}$ are bound to happen. We quantify this statistic as an absolute difference,
$\hat{D}_{j,n}^{t}=|S_{j,n}^{t}-\hat{S}_{j,n}^{t}|$ (6)
@@
We emphasize that since Eq. 5 leverages sensitivity from previously computed batches and domains, $\hat{D}_{j,n}^{t}$ would act as a good indicator of how uncertain $S_{j,n}^{t}$ is to the previous batches and domains i.e. $\hat{S}_{j,n}^{t}$, for $\theta_{j,n}^{t}$. This would take care of error accumulation too. To automatically alter the LRs of these parameters, the “importance" ($i_{j,n}^{t}$) is calculated as,
$i_{j,n}^{t}=\hat{D}_{j,n}^{t}/\hat{S}_{j,n}^{t}$ (7)
@@
Eq. 6 enforces $\hat{D}_{j,n}^{t}$$\geq$0 which, in turn, makes $i_{j,n}^{t}$$\geq$0. And so, the final learning rate to update $\theta_{j,n}^{t}$ would be,
$\hat{\kappa}_{j,n}^{t}=\kappa*i_{j,n}^{t}$ (8)
@@
where $\kappa$ is the base fixed LR. The intuition behind the formulation of Eq. 7 is - smaller values of $\hat{S}_{j,n}^{t}$ indicate lower sensitivity, and hence larger parameter updates are required via promoting the LR. At the same time, large uncertainties by $\hat{D}_{j,n}^{t}$ would also require larger updates. To avoid any overflow/underflow, we add $\epsilon$ (0$<$$\epsilon$$\ll$1) to the numerator and the denominator in Eq. 7.
@@ 4.3 Optimization Objective
Though our method is robust, it is necessary to avoid miscalibrated labels. TENT [36] minimized the entropy on the output logits, l, following $\mathcal{H}(l)=-\sum_{c}p(l_{c})\log(p(l_{c}))$, where $p(l_{c})$ is the probability for class c. Following EcoTTA [33], we use entropy minimization as the objective for adaptation where $\mathcal{H}(l)$ is minimized for samples below a certain fixed threshold $\mathcal{H}_{0}$ as,
$\mathcal{L}^{entropy}=\mathbf{1}_{\mathcal{H}(l)\leq\mathcal{H}_{0}}\mathcal{H}(l)$ (9)
$\mathcal{H}_{0}$ is set to $0.4\times\log C$ following EcoTTA. For a fair comparison with prior works [37, 24], we apply a consistency loss between l and $\hat{l}$, computed as the output logits from the augmented version of $x_{k}$, to regularize the model as follows,
$\mathcal{L}^{const}=-\sum_{c}p(l_{c})\log(p(\hat{l}_{c}))$ (10)
where, $p(\hat{l}_{c})$ is the probability for class c. Our overall optimization objective is as follows,
@@
where $\lambda$ is a regularisation coefficient. In summary, upon calculating the parameter-specific LRs (Eq. 8), we optimize the continually fine-tuned model to minimize Eq. 11.
@@ 5 Experiments
We draw a contrast of our proposed method against other state-of-the-art works in a continual test-time adaptation (CTTA) setting. However, to have a broader study and analysis, we also report results in a gradual test-time adaptation (GTTA) setting, following [37, 24]. In the following subsections, we describe the implementation details of both the challenging adaptation settings.
@@ 5.1 Experimental Settings and Baselines
Datasets and tasks. Following the standard benchmarks set by CoTTA [37], we evaluate our proposed method on CIFAR-10C, CIFAR-100C, and ImageNet-C, based on image corruption schemes as set in [9]. Each dataset contains a set of 15 corruption styles as tasks (e.g. gaussian noise, shot noise, impulse noise, $\ldots$) with 5 severity levels. We follow CoTTA and apply the same test-time augmentations, like color jitter, gaussian blur, gaussian noise, random flips, and random affine, to the input $x_{k}$. As the evaluation metric, we report the mean classification error.
CTTA implementation details. We evaluate our approach following prior state-of-the-art methods [36, 37, 3, 33, 28, 24]. To maintain fairness regarding the usage of pre-trained source models, we employ WideResNet-28 [42] for CIFAR-10C, ResNeXt-29 [39] for CIFAR-100C, and ResNet-50 [8] for ImageNet-C, all available on RobustBench [4]. We optimize using an Adam optimizer, setting base learning rates ($\kappa$) to 5e-4 for CIFAR-10C and CIFAR-100C, and 5e-5 for ImageNet-C. For balancing parameter sensitivities, we set $\alpha$ to 0.5, 0.9, and 0.5 respectively, with temperature coefficients T set to 50, 100, and 1000 respectively. We set $\eta$ to 1, 0.5, and 0.3, and $\lambda$ to 0.01 throughout. Batch sizes are set to 200, 200, and 64 for each dataset, following CoTTA, and results are only reported for the highest severity level of 5 for each task.
GTTA implementation details. In the standard CTTA setup, each task only receives images with the highest severity. However, in GTTA, following CoTTA, for each task, the severity levels are gradually changed as $\ldots$3$\rightarrow$2$\rightarrow$1 $change]{task}$ 1$\rightarrow$2$\rightarrow$3$\rightarrow$4$\rightarrow$5$\rightarrow$4$\rightarrow$3$\rightarrow$2$\rightarrow$1 $change]{task}$ 1$\rightarrow$2$\rightarrow$3$\ldots$. We adhere to the details specified in the CTTA setting.
Baselines. We compare against several CTTA baselines, including BN Stats Adapt (BN-1) [28], TENT Continual [36], CoTTA [37], EcoTTA [33], AdaContrast [3], DSS [38], and LAW [24]. We also contrast our work against surgical fine-tuning [15]. For EcoTTA, we report the results on CIFAR-10C and ImageNet-C only since they do not use ResNeXt-29 [39] for CIFAR-100C. Following the findings of surgical fine-tuning [15], we continually fine-tune only the first convolutional block of the respective models, denoted as “Surgical fine-tuning (continual)", using an SGD optimizer with fixed learning rates of 1e-3 for the CIFAR datasets and 5e-4 for ImageNet-C. For the reported results of LAW, we reproduce their results strictly following their official implementation [^1] https://github.com/junia3/LayerwiseTTA/tree/main with parameters as mentioned in their paper.
@@
25.8 18.1 22.7 12.3 25.3 13.1 10.7 13.5 13.1 12.2 8.5 11.8 17.9 12.0 15.4 15.5
Table 1: Mean errors (%) on CIFAR-10C - CTTA mean errors of the 15 corruptions (tasks) at a severity level of 5.
@@
37.3 32.5 34.9 26.2 35.3 27.5 24.6 28.8 29.1 34.1 23.5 26.9 31.2 26.6 34.1 30.1
Table 2: Mean errors (%) on CIFAR-100C - CTTA mean errors of the 15 corruptions (tasks) at a severity level of 5.
@@ 5.2 CTTA Classification Results
Results on CIFAR-10C and CIFAR-100C. We contrast our method against all the baselines as in Tables 1 and 2. The results reveal that our method performs the best across all training setups and baselines for both datasets. We achieve large improvements of 4.9% and 5.2% for CIFAR-10C and 5.3% and 30.8% for CIFAR-100C, over BN-1 and TENT Continual respectively. Concerning recent works of LAW and DSS, our method improves 1.7% and 0.5% on CIFAR-10C and 1.6% and 0.8% on CIFAR-100C, respectively. It is noteworthy that TENT Continual initially excels in the early tasks but experiences error accumulation with time, leading to an increase in mean error. In the study by CoTTA, the method’s reliance on full model updates renders it computationally expensive and unsuitable for deployment. An interesting observation is that with a drastic reduction in the number of trainable parameters, as in “Surgical fine-tuning (continual)", the mean errors are better than most CTTA benchmarks, if not the best. LAW uses the uncertain pseudo labels to calculate the domain shift approximated via the FIM. The minor error that is introduced for each batch begins to propagate across all the domains. In addition, the imbalance caused by the “domain-level" FIM and the current batch’s FIM, results in a higher error rate for the initial tasks. In our method, the computation of the model prediction uncertainty to automatically select layers is useful (Eq. 1). This affirms the findings in [15] that concentrated adaptation, for image corruptions, in the initial layers suffices.
Results on Imagenet-C In Table 3, we report the CTTA results on the ImageNet-C dataset. As shown, our method outperforms all the previous CTTA baselines and “Surgical fine-tuning (continual)". We observe an improvement of 0.9% and 4.5% over LAW and DSS respectively. In addition, we achieve 2.6%, 5.4%, and 2.5% improvements over CoTTA, AdaContrast, and TENT Continual, respectively.
@@
81.1 73.3 70.8 77.0 71.8 62.3 53.9 56.7 60.8 50.4 36.2 65.9 48.0 45.2 48.0 60.1
Table 3: Mean errors (%) on ImageNet-C - CTTA mean errors of the 15 corruptions (tasks) at a severity level of 5.
@@
CIFAR-10C 24.7 13.7 20.4 10.9 12.1 11.3 10.9 9.5
CIFAR-100C 33.6 29.9 74.8 26.3 33.0 25.3 27.2 26.1
ImageNet-C 58.4 48.3 46.4 38.8 66.3 41.4 40.5 38.5
Table 4: Mean errors (in %) on CIFAR-10C, CIFAR-100C, and ImageNet-C - We report the GTTA mean errors averaged across all the severity levels (1$\rightarrow$2$\rightarrow$3$\rightarrow$4$\rightarrow$5).